OpenStudy (anonymous):

SOMEONE HELP PLEASE...AM LITERALLY BEGGING My family pool is 6 feet deep, 20 feet wide and 20 feet long. My dog pool is 5ft wide, 5ft long and 1.5 feet deep. What relationship does the volume of the dogs' pool have with the volume of the original family pool? How does this compare to the relationship in the dimensions of the dogs' pool to the dimensions of the family pool? What is the difference in volume when the pool is filled to the top versus filled to 6 inches below the top? Was the amount of pool liner material representative of the lateral or surface area of a rectagular

OpenStudy (anonymous):

of a rectangular prism? why or why not Imagine Mrs. Noether asked you to add steps to the family pool. How would adding steps affect the volume of the water? How would adding steps affect the amount of pool liner material needed? Explain why each change would take place.

OpenStudy (anonymous):

@AccessDenied, @telltoamit can you guy's help plz?

OpenStudy (jim766):

start by finding the volume of each pool V = l*w*H

OpenStudy (anonymous):

how do you do that again?

OpenStudy (jim766):

family pool length 20ft width 20 ft Height or depth 6 feet Vol = L * W * H substitute the numbers in and solve

OpenStudy (anonymous):

so just do 20x20x6?

OpenStudy (jim766):

right

OpenStudy (anonymous):

that gives me 1600 and for the dog pool its 37.5, so what do i do next?

OpenStudy (jim766):

20*20*6 = 2400 try this one again 5*5*1.5 = 37.5

OpenStudy (anonymous):

ohhh my bad

OpenStudy (jim766):

1st question how do the volumes compare? 2400 / 37.5

OpenStudy (jim766):

opps they want to know how dog compares to people pool so 37.5/ 2400

OpenStudy (anonymous):

yea and i dont understand how am i gonna compare them

OpenStudy (jim766):

\[37.5\div2400\] is what I meant above

OpenStudy (anonymous):

i got .015625

OpenStudy (jim766):

can you make it a fraction?

OpenStudy (anonymous):

no, am not good with doing that :/

OpenStudy (jim766):

are you using a graphing calculator?

OpenStudy (anonymous):

yes i am

OpenStudy (jim766):

TI 83 or TI 84?

OpenStudy (anonymous):

84

OpenStudy (jim766):

does it still say .015625 on the screen?

OpenStudy (anonymous):

yes

OpenStudy (jim766):

push math enter enter note it says frac it will turn your decimal into a fraction, try it

OpenStudy (anonymous):

wooooow, its 1/64

OpenStudy (jim766):

yeah, cool right? what that means is the dog pool is 1/64 the size of the people pool now question 2 compare the dimensions len width depth dog 5 5 1.5 People 20 20 6 to compare them do what you did before, divide each..... for ex \[dog lenth \div people length\]

OpenStudy (anonymous):

wait what do i divide, like i know its those numbers but can you start me of with one like is it 5/ 20 or something?

OpenStudy (jim766):

yes whats 5/20 (oh and make it a fraction)

OpenStudy (anonymous):

it would be .25 which in a fraction is 1/4 and 1.5/6 is the same 1/4

OpenStudy (jim766):

yep .... the dog dimensions are 1/4 the people dimensions.. do you know what 1/4 * 1/4 * 1/4 is?

OpenStudy (jim766):

do .25 * .25 * .25 if you want

OpenStudy (anonymous):

its 15625

OpenStudy (jim766):

.25 * .25 * .25 = .015625 or 1/64 question 3 find the volumes with 6 less inch of water what will the new depths be?

OpenStudy (anonymous):

and how do i do that, am sorry i am just clueless

OpenStudy (jim766):

dog depth is 1.5ft what would it be if you took away 6 inches (how much of a foot is 6 inches)

OpenStudy (anonymous):

0.5 feet

OpenStudy (jim766):

so the dog vol would be v = 5 * 5 * 1 then do the people pool by taking .5 ft off its depth v = 20 * 20 * 5.5

OpenStudy (jim766):

what is the difference of the two vol ? people org vol was 2400 then sub whatever you get above dog pool was 37.5, then sub what every you get above

OpenStudy (anonymous):

wait is that an answer or your asking me

OpenStudy (anonymous):

oh for 5x5x1 i got 25 and 20x20x5.5 is 2200

OpenStudy (jim766):

so 37.5 - 25 = the diff of full pool and pool 6n less and 2400 - 2200 =

OpenStudy (anonymous):

12.5, 200 is that the ans for number three?